how do you multiply out (x+2)^3
Answer by Alan3354(69443)

how do you multiply out (x+2)^3
(x+2)^3 = (x+2)*(x+2)*(x+2)
=(x^2 + 4x + 4)*(x+2)
=x^3 + 2x^2 + 4x + 8